Design the concrete mix according to the following conditions:



Design Environment



Building frame



Required design strength = 27.6 MPa



Minimum dimension = 150 mm



Minimum space between rebar = 40 mm



Minimum cover over rebar = 40 mm



Maximum slump is required



Statistical data indicate a standard deviation of compressive strength of 2.1 MPa is



expected (more than 30 samples).



Only air entrainer is allowed.




Available Materials



Air entrainer: Manufacture specification 6.3 ml/1% air/100 kg cement. (Use maximum



or +2% air content)



Coarse aggregate: 19 mm nominal maximum size, river gravel (rounded)



Bulk oven-dry specific gravity = 2.55, Absorption = 0.3%



Oven-dry rodded density = 1761 kg/m³



Moisture content = 2%



Fine aggregate: Natural Sand



Bulk oven-dry specific gravity = 2.6

1
Expert's answer
2022-03-17T14:55:51-0400

OR BUILDING SLUMP IS 25-100



FOR 19MM AGGREGATE SIZE


WATER REQUIRED IS 168KG/CUM



FOR STRENGTH OF CONCRETE =27.6




CORRESPONDING WATER CEMENT RATIO IS 0.48


WATER=168


Cement required =168/(0.4)=247kg/cum



for fineness of =2.47


and for aggregates size=19mm




Aggregate required =0.65 cum


density of AGGREGATE =1761



wt of aggregates =0.65"\times" ×1761=1144kg/cum




wt of cement =247




wt of water =168



wt of aggregates =1144



density of water =1000kg/cum



vol of water =0.168 cum



density of cement =1400kg/cum




vol of cement =0.176


vol of AGGREGATE =0.65


vol of water=0.168


vol of cement =0.176


due to moisture and absorption 


reduce volume 


volume =0.63


vol of water =0.15


vol of cement =0.16




vol of fine =1-0.63-0.15-0.16=0.06 cum




vol of agg=0.63


vol of fine =0.06


vol of water =0.15


vol of cement =0.16


wt of aggregates =0.63×1761

=1109kg


wt of water =150 kg


wt of cement =210 kg


wt of fine =2.659×1000×0.06=159.6 kg



These are required for 1 cum of concrete